x+y=4 is the perpendicular bisector of AB where A=(3,-3). Find B.
Answer by josgarithmetic(39618)

Point A is on the line, -x+y=c, chosen this way based on familiarity with Standard Form equation for a line; therefore, you can use the coordinate values of point A to compute the value for c.

Once found c, you can to find the intersection point of -x+y=c and x+y=4. This intersection point will be the midpoint of segment AB. You can use Midpoint Formula to find the coordinates for point B.